Grade 9Mechanics

A ball is thrown forward from the top of a cliff with a velocity of 10ms. The height of the cliff above the ground is 45m. Calculate the distance of the ball to the hirizontal just before hitting the ground

Stephen
6 Years agoGrade 9

4 Answers

Saurabh Koranglekar
6 Years ago
Dear student

Time for fall = sqrt( 2*h/g) = sqrt ( 2*45/10)= 3 sec

Horizontal distance travelled by the ball= 3 * speed horizontal
= 3* 10 = 30 m

Regards
Vikas TU
6 Years ago
Dear student 
Horizontal velocity is irrelevant. Horizontal and vertical motion are independent of each other.
h = Vin*t + 1/2 gt^2 
h = 1/2 gt^2 
From here
t = sqrt(2h/g)
t = 3sec 
Horizontal distance travelled by the ball= 3 * speed horizontal= 3* 10 = 30 m
aswanth nayak
6 Years ago
Dear Student,
 
Horizontal and vertical motion are independent of each other.
h = Vin*t + 1/2 gt^2 
h = 1/2 gt^2 
From here
t = sqrt(2h/g)
t = 3sec 
Horizontal distance travelled by the ball= 3 * speed horizontal= 3* 10 = 30 m
Horizontal velocity is irrelevant.
 
Regards
Sandeep
6 Years ago
Time for fall = sqrt( 2*h/g) = sqrt ( 2*45/10)= 3 secHorizontal distance travelled by the ball= 3 * speed horizontal= 3* 10 = 30 m